OpenStudy (anonymous):

A bicyclist traveled 280 km at a certain speed. If the speed had been increased 5 km/h, the trip could have been made in 1 hour less time. Find the actual speed.

OpenStudy (anonymous):

in first case v=280/x where v is velocity,t is time taken in the second case v+5=280/(x-1) use these equations to get the value of x in terms of v..then use the value of x in any one of above equations,you will get the required speed

OpenStudy (anonymous):

is there some way I can solve this by using quadratic formulas or squaring something so I can factor it?

OpenStudy (anonymous):

you can use the quadratic formula

OpenStudy (anonymous):

okay and how would I set that up..? Like the one up above..?

OpenStudy (anonymous):

from first case vx=280 from 2nd case (v+5)(x-1)=280 so vx=(v+5)(x-1) solve this to get x in terms of v and do as i said above

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

you're welcome